# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v2

@@ Abstract.
Natural Language Processing (NLP) is revolutionising the way legal professionals and laypersons operate in the legal field. The considerable potential for NLP in the legal sector, especially in developing computational tools for various legal processes, has captured the interest of researchers for years. This survey follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses framework, reviewing 154 studies, with a final selection of 133 after manual filtering. It explores foundational concepts related to NLP in the legal domain, illustrating the unique aspects and challenges of processing legal texts, such as extensive document length, complex language, and limited open legal datasets. We provide an overview of NLP tasks specific to legal text, such as Legal Document Summarisation, legal Named Entity Recognition, Legal Question Answering, Legal Argument Mining, Legal Text Classification, and Legal Judgement Prediction. In the section on legal Language Models (LMs), we analyse both developed LMs and approaches for adapting general LMs to the legal domain. Additionally, we identify 16 Open Research Challenges, including bias in Artificial Intelligence applications, the need for more robust and interpretable models, and improving explainability to handle the complexities of legal language and reasoning.
@@ 1. Introduction
In recent years, advancements in Natural Language Processing (NLP) have significantly impacted the legal domain by simplifying complex tasks, such as Legal Document Summarisation (LDS), Legal Argument Mining (LAM) enhancing legal text comprehension for laypersons, and improving Legal Question Answering (LQA) and Legal Judgement Prediction (LJP) (Savelka et al., 2023; Pont et al., 2023; Janatian et al., 2023; Gesnouin et al., 2024; Chlapanis et al., 2024; Jiang et al., 2024; Li et al., 2022; Huang et al., 2024). These improvements are primarily attributed to advancements in Neural Network (NN) architectures, such as transformer models (Vaswani et al., 2017). NLP techniques now enable machines to generate text, answer legal questions, drafting a regulation, and simulate legal reasoning, which can revolutionise legal practices (Janatian et al., 2023). Applications, such as contract review (Graham et al., 2023; Mamooler et al., 2022; Maroudas et al., 2022; Tuggener et al., 2020) and case prediction (Niklaus et al., 2021; Vianna et al., 2023) have been automated to a large extent, speeding up processes, reducing human error, and cutting operational costs (Zhong et al., 2020b). Additionally, the use of NLP allows lawyers and legal professionals to reduce their workload, enhance efficiency, and minimise errors in decision-making processes (Savelka et al., 2023; Huang et al., 2024). Despite the rapid development of NLP, challenges remain in processing lengthy documents, understanding complex language, and navigating complicated document structures (Yuen Iu and Man-Yi Wong, 2023; Vianna et al., 2023; Sovrano et al., 2021; Nguyen et al., 2024; Jiang et al., 2024; Ganguly et al., 2023; Bhattacharya et al., 2019; Huang et al., 2024), yet the promise of Large Language Models (LLMs) enhances the efficiency, fairness, and precision of legal services (Huang et al., 2024; Fagan, 2024).
@@
Despite these advantages, the integration of NLP in the legal domain is not without challenges such as, biases and unfairness, and explainability issues (Tamkin et al., 2021; Sheng et al., 2019; Deroy and Maity, 2023). The use of Artificial Intelligence (AI) in legal applications must follow strict standards of accuracy, fairness, and transparency, given the potential impact on clients’ lives and rights.
@@
This survey article explores the current landscape of NLP applications within the legal domain. It discusses its potential benefits and the practical challenges it poses. NLP is a broad field covering a wide range of techniques for processing, analysing, and understanding human language. By examining the latest advancements and applications of NLP in law, this article provides a comprehensive overview of the field. Table 1 summarises the scope of the survey and categorise the research into several areas: LQA, LJP, Legal Text Classification (LTC), LDS, legal Named Entity Recognition (NER), LAM, legal corpora, and legal Language Models (LMs). Each category lists relevant projects and papers and shows the work being done in each sub-field. Notably, there is comparatively less research in NER and legal corpora, whereas LDS and LQA have seen extensive research activity, with a substantial number of datasets and research contributions. This summary provides an overview of how NLP techniques are applied to various challenges in the legal domain and offers insights into future directions of AI in legal practice.
@@
Document Summarisation Gelbart and Smith (1991a, b); Moens et al. (1999); Farzindar (2004); Polsley et al. (2016); Schraagen et al. (2022); Zhong and Litman (2022); Moro et al. (2023); Jain et al. (2024); Liu et al. (2024); Shen et al. (2022)
Text Classification Elnaggar et al. (2018); Wei et al. (2018); Lee and Lee (2019); Bambroo and Awasthi (2021); Song et al. (2022); Wang et al. (2022); Mamooler et al. (2022); Chalkidis et al. (2019b); Tuggener et al. (2020); Chalkidis et al. (2021); Papaloukas et al. (2021); Graham et al. (2023); Nguyen et al. (2021); Bhattacharya et al. (2023); Galassi et al. (2024); Grabmair et al. (2015)
@@
Question Answering Huang et al. (2020); Sovrano et al. (2020); Khazaeli et al. (2021); Zhong et al. (2020c); Askari et al. (2022); Zhang et al. (2023c); Louis et al. (2024); Sovrano et al. (2024); Yuan et al. (2023); Askari et al. (2024); Chen et al. (2023); Büttner and Habernal (2024); Sovrano et al. (2021)
Table 1. An overview of the research areas in legal NLP and the key publications in the survey.
This document is organised as follows. Firstly, In Section 2, we discuss previous surveys in this multidisciplinary domain. In Section 3, we provide a detailed overview of legal language and the basic principles of NLP as they apply to the legal domain. In Section 4, we briefly explain the research methodology of this work and how we extracted the resources. In sections 5-11, we explore various NLP tasks that are tailored for legal applications and show their unique requirements and the methodologies employed to address them. Additionally, we study the datasets available for training and evaluating legal NLP tasks, emphasising their characteristics and the implications they have for model performance. Following this, in Section 12, we investigate the development of LMs that has been specifically adapted for the legal field. Finally, in Section 13, we address the key challenges associated with deploying NLP technologies in legal settings, discussing both current issues and potential solutions. Since this survey contains many acronyms, Table 2 provides the list of acronyms and their meanings to make it easier to follow.
@@ 2. Related Work
Several studies have examined the use of NLP in the legal domain, each focusing on different aspects. To provide a comprehensive understanding of the existing research on integrating AI within the legal domain, we present an overview of recent literature reviews, as summarised in Table 3. Most survey papers on intelligent legal systems focus either on traditional NLP technologies for specific tasks, such as LJP and LDS or take a broader approach but still miss certain aspects. As illustrated in Table 3, there seems to be a gap in survey papers that thoroughly examine all facets of this multidisciplinary field. Our current work aims to bridge this gap by offering a comprehensive survey of all NLP tasks, existing datasets and corpora, and LMs in the legal domain. We use ✓ to indicate papers that study the most of the existing research on each subject in legal NLP. Papers that do not address a subject receive ✗, and those that partially cover specific subjects along with their datasets or LMs are marked with –.
Dias et al. (2022) discussed AI and NLP concepts and their applications in the legal domain. The study did not analyse legal datasets, specific NLP tasks, or legal LLMs. This work, in contrast, reviews NLP tasks in the legal domain, including LQA, LDS, and LTC, along with legal datasets and corpora. Sun (2023) examined two NLP tasks in the legal domain, such as LJP and statutory reasoning. The study reviewed three datasets and two LLMs related to these tasks. Unlike Sun’s work, this survey covers a wider range of NLP tasks and includes legal corpora and datasets. Cui et al. (2023) focused on LJP, reviewing 43 datasets in nine languages. The study evaluated classification, text generation, and regression tasks. It also discussed pre-trained LMs used for LJP. This work differs by covering multiple NLP tasks, legal corpora, and dataset availability.
Anh et al. (2023) explored challenges in legal language processing and how LLMs can address them. The study summarised six NLP tasks and discussed ethical concerns, including bias, privacy, and transparency. While these issues are relevant, this survey focuses on existing NLP methods, datasets, and legal corpora for pre-training and fine-tuning.
Ganguly et al. (2023) surveyed legal text processing challenges, such as NER and sentence boundary detection. The study reviewed historical developments in AI and law research, as well as recent NLP advancements. It covered LDS and LJP but did not examine LQA, LTC, or legal corpora. Our work covers all these areas.
Chen et al. (2024a) studied LLMs in finance, healthcare, and law. While attempting to provide a broad view of LLM applications in the legal domain, the study’s scope resulted in a less detailed review of specific NLP tasks and datasets. Additionally, it did not examine large legal corpora or pre-training methods, aspects that our survey addresses.
@@
Table 3. Overview of existing surveys on NLP in the legal domain.
@@
The main difference between this work and previous surveys is that this survey aims to provide a more general description of all aspects of NLP tasks in the legal domain, rather than focusing solely on specific applications. The main contributions of this survey are summarised as follows: (1) This article extends beyond previous surveys by examining a broad spectrum of studies, and applications of legal NLP. By discussing datasets and large corpora in 24 languages and exploring the popular legal LMs, this survey establishes itself as an important resource in the field of legal NLP. (2) The survey offers an in-depth look at the challenges of integrating NLP with legal applications, with detailed discussions on technical solutions that tackle these issues, thereby enhancing understanding and encouraging further research in this evolving field. (3) This survey also highlights the existing research gaps in legal NLP, identifying areas that require further exploration and development, and providing a road-map for future research efforts in the legal NLP domain.
@@ 3. Background and Foundational Concepts
In this section, we will establish an essential understanding of how NLP intersects with the legal domain. Initially, we explore the characteristics of legal documents, which are fundamental to this intersection. Legal texts are known for their complex structure, specialised words. Recognising these attributes is important because they influence the development and use of AI in legal practices.
@@ 3.1. Legal Documents
Legal documents are typically written in a descriptive language and presented in an unstructured text format. They have unique features that set them apart from other fields. These documents cover a broad array of texts essential to the functioning of the legal system. They include court filings, judgements, legislation, treaties, contracts, and legal correspondence, each serving a specific purpose and following to specific formatting and content standards that reflect legal logic and hierarchy. Legal documents are Fundamental tools for lawyers, judges, and legal scholars, facilitating case analysis, legislative review, and contract drafting. They are also essential in legal education and practice and provide the basis for legal arguments and decisions. Common examples of legal documents include case law repositories, statutory databases, and collections of legal agreements. These documents are utilised in various legal processes such as drafting legal arguments, performing legal analysis, and ensuring regulatory compliance.
@@ 3.1.1. Legal language and its characteristics
Legal language is characterised by its unique features that set it apart from everyday language, primarily due to its function in the legal system. One prominent feature is its formality, where legal language often employs a more formal vocabulary and syntax to ensure precision and avoid ambiguity (Gibbons and Turell, 2008). This formality is important, as the meaning of terms can have legal effects. Legal texts also typically utilise passive constructions and complex sentence structures to provide detailed and comprehensive descriptions (Gibbons and Turell, 2008). These constructions help clarify responsibilities and outcomes without directly attributing actions or intentions to specific parties.
@@
Another distinctive aspect of legal language is its reliance on specialised words and phrases. This includes terms that have specific meanings within legal contexts, archaic words that are not commonly used in everyday language, and standardised phrases that have been historically embedded in legal tradition (Gibbons and Turell, 2008). This can make legal documents less accessible to non-specialists, requiring legal professionals to interpret the content accurately.
@@
Furthermore, legal language is heavily intertextual, meaning it frequently references other legal texts, such as statutes, regulations, and case law. This characteristic ensures that legal arguments are grounded in and supported by existing legal frameworks and previous cases. The dense use of citations and references in legal documents not only supports the arguments made but also connects the document to a broader legal discourse. Such intertextuality demands that legal professionals not only understand the texts themselves but also the broader legal context in which they operate. To illustrate the intertextuality of legal language, Figure 1 shows a sample page from the Code of Federal Regulations of the US, extracted from the Electronic Code of Federal Regulations [^1] https://www.ecfr.gov. Figure 1 displays § 40.51, labour Certification, from 22 C.F.R. § 40.51, which is part of the Code of Federal Regulations of the US. This section falls under Title 22, governing regulations related to foreign relations, specifically detailing the requirements and procedures for labour certification. Notably, the text includes highlighted references to other legal sources, such as INA 212(a)(5). This citation refers to the Immigration and Nationality Act, specifically section 212, which outlines various conditions for inadmissibility into the US, under subsection (a), paragraph (5). In the “Credits” part, you can see the reference “56 FR 30422, July 2, 1991,” which points to a publication in the Federal Register. Here, “56 FR” indicates the volume number, and “30422” is the page number where the document begins. The date “July 2, 1991,” marks the publication date in the Federal Register.
@@
Figure 1. A sample page from the Code of Federal Regulations, illustrating the structured and referenced nature of legal documents.
@@
A sample page from the Code of Federal Regulations of the US
The specific characteristics of legal language, such as its formal vocabulary, complex syntax, and extensive use of references, present many challenges for NLP. Disambiguation titles and nested entities are other issues in legal contexts (Krasadakis et al., 2024). Disambiguation titles, such as ‘The President of USA’ requires precise identification based on contextual details, such as time and location. Nested entities, where titles of legislative articles refer to other laws, introduce further complexity. On the other hand, legal documents are frequently provided in non-machine-readable PDF formats, complicating data extraction and processing. Additionally, the variation in legal reasoning, which includes rule-based, analogical, and evidentiary arguments, along with changes in legal standards, creates challenges for applying conventional NLP models (Freeman Engstrom and B. Gelbach, 2021). Further elements, such as the peculiar use of punctuation affecting text segmentation and the frequent presence of digits and numbers, can also disrupt traditional NLP pipelines. These challenges show the need for specialised NLP solutions tailored to the legal domain.

### v3

# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models and Challenges

@@ Abstract.
Natural Language Processing (NLP) is revolutionising the way both professionals and laypersons operate in the legal field. The considerable potential for NLP in the legal sector, especially in developing computational assistance tools for various legal processes, has captured the interest of researchers for years. This survey follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses framework, reviewing 154 studies, with a final selection of 131 after manual filtering. It explores foundational concepts related to NLP in the legal domain, illustrating the unique aspects and challenges of processing legal texts, such as extensive document lengths, complex language, and limited open legal datasets. We provide an overview of NLP tasks specific to legal text, such as Document Summarisation, Named Entity Recognition, Question Answering, Argument Mining, Text Classification, and Judgement Prediction. Furthermore, we analyse both developed legal-oriented language models, and approaches for adapting general-purpose language models to the legal domain. Additionally, we identify sixteen open research challenges, including the detection and mitigation of bias in artificial intelligence applications, the need for more robust and interpretable models, and improving explainability to handle the complexities of legal language and reasoning.
@@ 1. Introduction
Advancements in Natural Language Processing (NLP) have significantly impacted the legal domain by simplifying complex tasks, such as Legal Document Summarisation (LDS), Legal Argument Mining (LAM), enhancing legal text comprehension for laypersons, and improving Legal Question Answering (LQA) and Legal Judgement Prediction (LJP) (Savelka et al., 2023; Pont et al., 2023; Janatian et al., 2023; Gesnouin et al., 2024; Chlapanis et al., 2024; Jiang et al., 2024; Li et al., 2022; Huang et al., 2024). These improvements – like in many other data-driven fields – are primarily attributed to advancements in Neural Network (NN) architectures, such as transformer models (Vaswani et al., 2017). NLP techniques now enable machines to generate text, answer legal questions, draft regulations, and simulate legal reasoning, which have the potential to revolutionise legal practices (Janatian et al., 2023). Applications, such as contract review (Graham et al., 2023; Mamooler et al., 2022; Maroudas et al., 2022; Tuggener et al., 2020) and case prediction (Niklaus et al., 2021; Vianna et al., 2023) have been automated to a large extent, speeding up processes, reducing human error, and cutting operational costs (Zhong et al., 2020b). Additionally, the use of NLP allows lawyers and legal professionals to reduce their workload, enhance their efficiency, and minimise errors in decision-making processes (Savelka et al., 2023; Huang et al., 2024). Despite rapid developments in NLP technology, significant challenges remain in the legal context due to lengthy documents, complex legal language, and complicated document structures (Yuen Iu and Man-Yi Wong, 2023; Vianna et al., 2023; Sovrano et al., 2021; Nguyen et al., 2024; Jiang et al., 2024; Ganguly et al., 2023; Bhattacharya et al., 2019; Huang et al., 2024).
@@
Despite these advantages, the integration of NLP in the legal domain is not without challenges, especially in terms of fairness, bias, and explainability issues (Tamkin et al., 2021; Sheng et al., 2019; Deroy and Maity, 2023). The use of Artificial Intelligence (AI) in legal applications must follow strict standards of accuracy, fairness, and transparency, given the potential impact on clients’ lives and rights. Nonetheless, Large Language Models (LLMs) have demonstrated potential to enhance the efficiency, fairness, and precision of legal tasks (Huang et al., 2024; Fagan, 2024).
@@
This survey article explores the current landscape of NLP applications within the legal domain. It discusses its potential benefits and the practical challenges it poses. NLP is a broad field covering a wide range of techniques for processing, analysing, and understanding human language. By examining the latest advancements and applications of NLP in law, this article provides a comprehensive overview of the field. Table 1 summarises the scope of the survey and categorise the research into several areas: LQA, LJP, Legal Text Classification (LTC), LDS, legal Named Entity Recognition (NER), LAM, legal corpora and legal Language Models (LMs). Each category lists relevant projects and papers, and shows the work being done in each sub-field. Notably, there is comparatively less research in NER and legal corpora, whereas LDS and LQA have seen extensive research activity, with a substantial number of datasets and research contributions. This summary provides an overview of how NLP techniques are applied to various challenges in the legal domain and offers insights into future directions of AI in legal practice.
Table 1. An overview of the research areas in legal NLP and the key publications discussed in this survey.
@@
Document Summarisation Gelbart and Smith (1991a, b); Moens et al. (1999); Farzindar and Lapalme (2004); Polsley et al. (2016); Schraagen et al. (2022); Zhong and Litman (2022); Moro et al. (2023); Jain et al. (2024); Liu et al. (2024); Shen et al. (2022)
Text Classification Elnaggar et al. (2018); Lee and Lee (2019); Bambroo and Awasthi (2021); Song et al. (2022); Wang et al. (2022); Mamooler et al. (2022); Chalkidis et al. (2019b); Tuggener et al. (2020); Chalkidis et al. (2021); Papaloukas et al. (2021); Graham et al. (2023); Nguyen et al. (2021); Bhattacharya et al. (2023); Galassi et al. (2024); Grabmair et al. (2015)
@@
Question Answering Huang et al. (2020); Khazaeli et al. (2021); Zhong et al. (2020c); Askari et al. (2022); Zhang et al. (2023c); Louis et al. (2024); Sovrano et al. (2024); Yuan et al. (2023); Askari et al. (2024); Chen et al. (2023); Büttner and Habernal (2024); Sovrano et al. (2021)
This document is organised as follows. Firstly, In Section 2, we discuss previous surveys in this multidisciplinary domain. In Section 3, we provide a detailed overview of legal language and the basic principles of NLP as they apply to the legal domain. In Section 4, we briefly explain the research methodology of this work and how we extracted the resources. In sections 5-11, we explore various NLP tasks that are tailored for legal applications, focusing on their unique requirements and the methodologies employed to address them. Additionally, we study the datasets available for training and evaluating legal NLP tasks, emphasising their characteristics and the implications they have for model performance. Following this, in Section 12, we investigate the development of LMs that have been specifically adapted to the legal field. Finally, in Section 13, we address the key challenges associated with deploying NLP technologies in legal settings, discussing both current issues and potential solutions. Since this survey contains many acronyms, Table 2 provides the list of acronyms and their meanings to make it easier to follow.
@@ 2. Related Work
Several studies have examined the use of NLP in the legal domain, each focusing on different problems and applications. To provide a comprehensive understanding of the existing research on integrating AI within the legal domain, we present an overview of recent literature reviews, as summarised in Table 3. Most survey papers on intelligent legal systems focus either on traditional NLP technologies for specific tasks, such as LJP and LDS, or adopt a broader approach but still overlook certain applications. As illustrated in Table 3, there is yet to be a comprehensive survey that thoroughly examine all facets of this multidisciplinary field; Our current work aims to bridge this gap by offering a comprehensive survey of all NLP tasks, existing datasets and corpora and LMs in the legal domain.
Dias et al. (2022) discussed AI and NLP concepts and their applications in the legal domain. Their study did not analyse legal datasets, specific NLP tasks, or legal LLMs. Our work, in contrast, reviews NLP tasks in the legal domain, including LQA, LDS and LTC, along with legal datasets and corpora. Sun (2023) examined two NLP tasks in the legal domain, such as LJP and statutory reasoning. Their study reviewed three datasets and two LLMs related to these tasks. Unlike Sun’s work, our survey covers a wider range of NLP tasks and includes legal corpora and datasets. Cui et al. (2023) focused on LJP, reviewing 43 datasets in nine languages. Their study evaluated classification, text generation and regression tasks. It also discussed pre-trained LMs used for LJP. Our work differs by covering multiple NLP tasks, legal corpora and dataset availability.
Anh et al. (2023) explored challenges in legal language processing and how LLMs can address them. Their study summarised six NLP tasks and discussed ethical concerns, including bias, privacy and transparency. While these issues are relevant, our survey focuses on existing NLP methods, datasets and legal corpora for pre-training and fine-tuning; we discuss elements of bias, fairness, privacy, interpretability, and explainability in Section 13.
Ganguly et al. (2023) surveyed legal text processing challenges, such as NER and sentence boundary detection. Their study reviewed historical developments in AI and law research, as well as recent NLP advancements. It covered LDS and LJP but did not examine LQA, LTC or legal corpora. Our work covers all of these areas.
Chen et al. (2024a) studied LLMs in finance, healthcare and law. While attempting to provide a broad view of LLM applications in the legal domain, their study’s scope resulted in a less detailed review of specific NLP tasks and datasets. Additionally, it did not examine large legal corpora or pre-training methods, aspects that our survey addresses.
@@
Table 3. An overview of existing surveys on NLP in the legal domain. We use a check mark (✓) to indicate papers that study the most of the existing research on each subject in legal NLP. Papers that do not address a subject receive a cross (✗), and those that partially cover specific subjects are marked with a dash (–).
@@
The main difference between our work and previous surveys is that our survey aims to provide a more general view of all aspects of NLP tasks in the legal domain, rather than focusing solely on specific applications. The main contributions of this survey are summarised as follows: (1) This article extends previous surveys by examining a broad spectrum of studies and applications of legal NLP. By discussing datasets and large corpora in 24 languages and exploring popular legal LMs, this survey establishes itself as an important resource in the field of legal NLP. (2) The survey offers an in-depth look at the challenges of integrating NLP with legal applications, with detailed discussions on technical solutions that tackle these issues, thereby enhancing understanding and encouraging further research in this evolving field. (3) This survey also highlights the existing research gaps in legal NLP, identifying areas that require further exploration and development and providing a road-map for future research efforts in the legal NLP domain.
@@ 3. Background and Foundational Concepts
In this section, we explain how NLP can be applied within the legal domain. We begin by outlining the unique characteristics of legal documents and legal language, highlighting the challenges these features pose for NLP applications. We then introduce core NLP concepts, methods, and paradigms relevant to legal texts, before discussing the recent impact of LLMs in the legal sector. Finally, we summarise key journals, conferences, and workshops that shape the field of legal NLP.
@@ 3.1. Legal Documents
Legal documents, such as court filings, judicial judgements, statutes, treaties, contracts, and formal legal correspondence encode the authoritative rules, rights, and duties that underpin our legal systems. They transmit legally operative information, including procedural requirements, enforceable obligations, and interpretative reasoning. Lawyers, judges, regulators, and academics consult these sources to analyse cases, interpret legislation, draft or negotiate agreements, verify compliance, and support teaching. Access is typically provided through official court portals, statutory databases, and commercial research platforms.
@@ 3.1.1. Legal language and its characteristics
Legal language is characterised by unique features that distinguish it from everyday language, primarily because of its role within the legal system. One prominent feature is its formality, where legal texts often employ a more formal vocabulary and syntax to ensure precision and avoid ambiguity (Gibbons and Turell, 2008). This formality is important, as the meaning of terms can have legal effects. Specialised vocabularies, fixed syntactic patterns, and even punctuation are chosen to maximise precision and avoid ambiguity, as small drafting choices can decisively alter legal consequences. One such example appeared in the 2006 dispute between Rogers Communications and Bell Aliant, [^1] https://crtc.gc.ca/eng/archive/2007/dt2007-75.htm where a single comma in the English version of a termination clause was interpreted to permit early cancellation of a multimillion-dollar pole-access contract, whereas the absence of that comma in the equally-authentic French text preserved the original five-year lock-in. This case illustrates how punctuation alone can shift rights and liabilities. Legal documents also typically use passive constructions and complex sentence structures to provide detailed and comprehensive descriptions (Gibbons and Turell, 2008) without directly attributing actions or intentions to specific parties.
@@
Another distinctive aspect of legal language is its reliance on specialised words and phrases. These include terms with specific legal meanings, archaic words rarely used in everyday language, and standardised phrases embedded in legal tradition (Gibbons and Turell, 2008). Such language can make legal documents less accessible to non-specialists, necessitating accurate interpretation by legal professionals.
@@
Furthermore, legal language is heavily intertextual, frequently referencing other legal texts, such as statutes, regulations, and case law. This ensures that legal arguments are grounded in existing legal frameworks and previous cases. The extensive use of citations and references situates each document within a wider legal discourse. Such intertextuality requires legal professionals to understand both the texts themselves and the broader legal context in which they operate. To illustrate the intertextuality of legal language, Figure 1 shows a sample page from the Code of Federal Regulations (CFR) of the US, extracted from the Electronic CFR [^2] https://www.ecfr.gov, which displays § 40.51, Labour Certification, of 22 CFR. This section is part of Title 22 of the CFR, which governs foreign relations and specifically details the requirements and procedures for labour certification. The text includes underlined references to other legal sources, such as INA 212(a)(5). This citation refers to section 212 of the Immigration and Nationality Act, subsection (a), paragraph (5), which outlines conditions for inadmissibility to the US. In the “Source” section, the citation “56 FR 30422, July 2, 1991,” indicates a Federal Register publication: “56 FR” is the volume number, “30422” the page where the document begins and “July 2, 1991” the publication date.
@@
Figure 1. A sample page from the CFR, illustrating the structured and referenced nature of legal documents.
@@
A sample page from the CFR of the US
Disambiguation titles and nested entities are other issues in legal contexts (Krasadakis et al., 2024). Disambiguation titles, such as “The President of USA” require precise identification based on contextual details, such as time and location. Nested entities, where titles of legislative articles refer to other laws, introduce further complexity. To further complicate matters, legal documents are frequently provided in non-machine-readable PDF formats, complicating data extraction and processing. Additionally, the variation in legal reasoning, which includes rule-based, analogical and evidentiary arguments, along with changes in legal standards, creates challenges for applying conventional NLP models (Freeman Engstrom and B. Gelbach, 2021). Elements such as the peculiar use of punctuation affecting text segmentation and the frequent presence of digits and numbers, can also disrupt traditional NLP pipelines. These challenges demonstrate the need for specialised NLP solutions tailored to the legal domain.